# Explanation Evaluation

### v1

@@ Results.
Figure 5 shows Spearman’s rank correlation with the original gold ratings obtained by aggregating the labels of all five raters. Humans only denotes the correlation between human majority-voted labels only, and others when including the respective LLM as an additional rater.
Figure 5: Change in Spearman’s rank correlation when using fewer raters and when an LLM is added as an additional rater. E.g., 4H+LLM compares the correlation between the original (five-rater) gold labels, majority-voted labels with four human raters and the respective LLMs added as a fifth rater, and four human raters only.
The leftmost cluster compares majority-voted labels between four human raters (4H) and four raters with an LLM added (4H+LLM). All correlations stayed the same (0.91), indicating that the LLMs did not compensate for the difference caused by a missing human rater. In this case, there was no harm nor benefit to including an LLM.
In the center, we observe the effect when the human raters are reduced to three (3H), with the influence of the LLM becoming stronger (3H+LLM). Human-only correlation with the original gold labels dropped to 0.86, and all LLMs decreased the correlation further. In this case, including an LLM was detrimental, bringing the majority labels further from the original human consensus.
Finally, the rightmost cluster shows the case of two human raters (2H) and adding the respective LLMs (2H+LLM). This represents the minimum possible ratio of humans to LLMs where humans still retain a majority. Interestingly, while overall correlation yet again decreased, GPT-4 improved the correlation (0.81 to 0.83).
Overall, we conclude that when there is already a large pool of workers (four+), there is no benefit to adding an LLM as a judge. GPT-4 can be a useful addition if there are only two human raters, but the resulting labels are still less reliable than those from three or more human raters.
@@ 4 Discussion
From our results, we conclude that most models we tested are not suitable for replacing human raters based on their divergence from human rater’s distributions. The best-performing model, GPT-4, generated labels that were particularly correlated (>0.90) with human majorities in supports and contrastive criteria but less so in others (0.52-0.78). Acting as an additional rater, GPT-4 did not affect the outcome over having one less human rater, harmed the outcome when there were three human raters, and improved the outcome when there were only two human raters. This suggests that GPT-4 may only be useful as an additional rater when human raters are particularly scarce.
From this, we concluded that LLMs are not a reliable replacement for human raters, except in isolated cases. However, in this work, we operated under the assumption that majority-voted labels are the ground truth. In practice, this may not always be the case, and the reliability of LLMs may vary depending on what information is desired in a given scenario. This opens up a new avenue for future work, where LLMs can be applied in a more nuanced manner, considering the context of the task and the intended use of the labels.
Overall, as a balanced approach, we recommend using LLMs in configurations that end with targeted human involvement, such as extensively using LLMs in development or filtering stages and experts for final testing or evaluation.
@@ 5 Related Work
LLMs as data labelers, and more broadly humans-and-LLMs-in-the-loop settings, are an emerging direction in data collection and labeling. E.g., Wiegreffe et al. (2022) developed a predict-and-explain pipeline that combines GPT-3 with a supervised filter trained on binary acceptability judgments from humans. More recently, Chiang and Lee (2023) proposed using LLMs to evaluate text, closely in line with our work. They, however, found the models successful in their setting. In contrast, we focus on explanation rating with more complex fine-grained criteria, and closely scrutinize potential weaknesses; Previous works brought into question the reliability of LLMs’ predictions, especially in prompting setups. E.g., Webson and Pavlick (2022) found that instruction-tuned models often produced good predictions with irrelevant or misleading prompts, bringing into question their real “understanding” of the task. Others reported unreliability of LLMs as labelers in various settings (Albrecht et al., 2022; Shen et al., 2023; Hada et al., 2023, i.a.), a line of work which we join with findings in the context of explanation evaluation.
@@
For explanations in the form of textual justifications, previous works often defined their own criteria for evaluation. Automatic evaluation borrowed from machine learning and measures overlap with “gold standard” text using (a) word-overlap metrics, e.g., BLEU, METEOR and ROUGE; and (b) embedding-based metrics, e.g., BERTScore and BLEURT (Clinciu et al., 2021). In contrast, human-tagged measures are more diverse and explanation-specific. For example, Clinciu et al. (2021) measured Informativeness and Clarity; Wiegreffe et al. (2022), inspired by social sciences, measured Acceptability, Generality, Factuality, Grammar, New Info, Supports Label, and Amount of Information; while Aggarwal et al. (2021) defined the criteria of Refutation, Complete, Comprehensive, Minimal, and Coherent. A recent study instead focused on the utility of explanations, i.e., their helpfulness in answering a question from a human point of view (Joshi et al., 2023). In our work, we largely followed these existing criteria, with the addition of a “supports” criterion indicating which answer the explanation attempts to justify.
@@
Explanations are often generated in predict-and-explain settings, where models provide justifications for their answers in QA-based benchmarks (e.g., Clinciu et al., 2021), or an elaborate-then-predict setting, where models output a prediction guided by its intermediate outputs such as knowledge statements or reasoning chains (e.g., Marasovic et al., 2022; Wang et al., 2023). In this work, we evaluate the capabilities of LLMs to evaluate explanations in an explain-and-predict setting, where models provide justifications for their answers in commonsense reasoning benchmarks. The following sections provide background on each aspect.
@@ 6 Conclusions
Using a newly-built dataset of free-text explanations and crowdsourced aspect-wise quality ratings, we analyzed the viability of LLMs as explanation evaluators. We found that LLMs decrease inter-annotator agreement when replacing human ratings, and that their ratings are moderately to highly correlated with human ratings, depending on the quality aspect. In a scenario where LLMs are used as additional raters instead, we found that they can improve the outcome when there are only two human raters, but are neutral to detrimental when there are three or more human raters. We conclude that while LLMs can provide ratings moderately consistent with an average crowdworker, they are not yet reliable enough for complete replacement.
@@
Some of the datasets used in our research have been found to contain biases that can be further perpetuated through explanations supporting them. As such, this data should not be used in any downstream applications without first controlling for potentially harmful samples. Some of the commercial LLMs we used have built-in filters for potentially harmful content, which were triggered during several of our experiments.
@@ Appendix A Dataset Statistics
A.1 Data Sources
Table 2 shows a breakdown of the samples in the test set per source dataset as described in Section 2.2.
@@ A.1 Data Sources
A.2 Data Examples
Table 4 and Table 5 show examples of best-rated and worst-rated explanations from each source dataset.
A.3 Label Distributions
-1 0 1 2 3 4 5
Supp. 11% 32% 31% 8% 9% 8%
Ovr. 11% 14% 30% 35% 10%
Well-wr. 28% 72%
Rel. 7% 93%
Fact. 12% 4% 85%
New i. 19% 46% 33% 2%
Unn. i. 17% 83%
Contr. 58% 42%
Table 3: Label distributions per criterion with human ratings. -1 denotes "none" for supports and "N/A" for factual. Scalar aspects are rounded to integer values.
Table 3 shows the label distributions of each criterion in our test set. Scalar aspects (overall rating, new information) are bucketed. ‘-1’ labels represent None for supports and N/A for factual.
@@ Appendix C Preliminary Experiment: Prompting Strategies
We compared single and compound calling, where the former prompts the model for a single criterion at a time, and the latter prompts the model for all criteria at once. We also compared default, averaged, and verbatim prompt formats, corresponding to a simple prompt with the explanation and the rating criteria, a voting mechanism over several prompt variants (described in Appendix D), and input identical to the human annotation guidelines, respectively. For single verbatim calls, only the relevant sections (guidelines and examples) for the target criterion were included. Default and averaged prompts were further compared in zero-shot and three-shot settings, where the latter contained the same examples as shown in the human guidelines.
All models were most successful with verbatim compound prompting, corresponding to the same guidelines as those shown to humans. Unless otherwise stated, all results presented in the main text are based on this setting.
Appendix D Prompt Variant Examples
This section shows the prompt variants used for prompt-averaged settings. Parts identical to the simple prompt are omitted and marked with ‘...’ for brevity. In few-shot settings, the initial guidelines are only shown once, followed by $N$ examples with a simple corresponding answer format, unless otherwise specified. E.g., for the simple prompt:
{mdframed}
[linewidth=0.5pt, linecolor=black]
QUESTION: <question>
a) <answer choice>
b) <answer choice>
EXPLANATION: <explanation>
1. a
2. 2
3. no
4. yes
5. yes
6. sufficient
7. no
8. yes
D.1 Simple Prompt
A straightforward prompt outlining the evaluation criteria.
Evaluate the given explanation according to
    the following criteria:
1. Which answer does it support? (a, b, c, d, e, none)
2. Overall rating? (1, 2, 3, 4, 5)
3. Is it well-written? (no, yes)
4. Is the explanation related to the question and answers? (no, yes)
5. Are all contained facts correct? (N/A, no, yes)
6. How much new information is given? (none, some, sufficient, ample)
7. Any unnecessary information? (no, yes)
8. Is it contrastive? (no, yes)
QUESTION: <question>
a) <answer choice>
b) <answer choice>
EXPLANATION: <explanation>
D.2 Shuffled Criteria Prompt
This prompt uses the same wording as the simple prompt but asks for ratings in a shuffled order.
Evaluate the given ...
1. How much new information is given? (none, some, sufficient, ample)
2. Is it contrastive? (no, yes)
3. Is the explanation related to the question and answers? (no, yes)
4. Is it well-written? (no, yes)
5. Overall rating? (1, 2, 3, 4, 5)
6. Which answer does it support? (a, b, c, d, e, none)
7. Are all contained facts correct? (N/A, no, yes)
8. Any unnecessary information? (no, yes)
QUESTION: ...
D.3 Slightly Paraphrased Prompt
This prompt slightly differs from the simple prompt in wording choice.
Evaluate the given explanation based on these criteria:
1. Which answer does it support? (a, b, c, d, e, none)
2. Rate the overall explanation (1, 2, 3, 4, 5)
3. Is it well-written? (no, yes)
4. Is the explanation relevant to the question and answers?  (no, yes)
5. Are all provided facts accurate? (N/A, no, yes)
6. How much new information is given? (none, some, sufficient, ample)
7. Does it contain unnecessary information? (no, yes)
8. Is it contrastive? (no, yes)
QUESTION: ...
D.4 Verbose Prompt
Here, the rating questions are more elaborated, but without adding more information compared to the other prompts.
Evaluate the explanation provided based on the following criteria:
1. Which answer choice does the explanation support? (a, b, c, d, e, none)
2. Rate the overall quality of the explanation. (1, 2, 3, 4, 5)
3. Is the explanation well-written? (no, yes)
4. Is the explanation related to the question and answer choices? (no, yes)
5. Verify the accuracy of all facts, if any, mentioned in the explanation. (N/A, no, yes)
6. How much new information is given in the explanation? (none, some, sufficient, ample)
7. Does the explanation contain any unnecessary information? (no, yes)
8. Is the explanation contrastive? (no, yes)
QUESTION: ...
D.5 Repeated Instructions Prompt
This prompt is identical to the simple prompt except for the example formatting. Instead of showing simple successions of questions, explanations, and ratings, the full instructions are repeated every time.
D.6 Swapped Instructions Prompt
The wording is the same as the simple prompt, but the order in which the instructions are presented is swapped. I.e., the example is shown first, then the general instruction text, followed by the rating questions. Similarly to the repeated instructions prompt (§D.5), the full instructions are repeated every time.
QUESTION: ...
EXPLANATION: ...
Evaluate the given explanation according to the following criteria:
1. Which answer does ...
Appendix E Crowdsourced Explanations for BCOPA
@@
The workers were explicitly instructed to write contrastive explanations which “focus on what the key difference between the options is, and how it leads to it being the correct choice in one case and not in the other.” This instruction was inspired by insights from Miller (2017). Figure 6 shows the crowdsourcing form. We collected two explanations per BCOPA question, totaling 3,000 explanations. Out of these, a random sample of 500 explanations was rated and included in our test set. Example explanations can be seen in Appendix A.2.
@@
CoS-E (best) Which effect of stress could cause death if not treated immediately? a) age grieving person b) heart attacks c) depression d) hair loss e) headache Explanation: heart attacks can stop your heart from properly pumping blood, thus leading to death. Overall: 4.4 Well-written: Yes
Supports: b) Related: Yes
Factual: Yes
New info: 1.8
Unn. info: No
Contrastive: No
@@
CoS-E (worst) The college kids weren’t known for their apartment decor choices, they used a folding chair inside their what? a) bathroom b) closet c) school d) room e) church Explanation: after years of being held captive in an isolated Overall: 1.0 Well-written: No
Supports: none Related: No
Factual: N/A
New info: 0.2
Unn. info: Yes
Contrastive: No
@@
CoS-E + fluency fixed (best) What kind of driving leads to accidents? a) stressful b) dangerous c) fun d) illegal e) deadly Explanation: Driving dangerously creates accidents. Overall: 4.0 Well-written: Yes
Supports: b) Related: Yes
Factual: Yes
New info: 0.6
Unn. info: No
Contrastive: No
@@
CoS-E + fluency fixed (worst) What is a grumpy person likely to have? a) headache b) understand themselves c) compliments d) own house e) happiness Explanation: A peculiar person. Overall: 1.0 Well-written: No
Supports: None Related: No
Factual: N/A
New info: 0.2
Unn. info: Yes
Contrastive: No
@@
Generated (best) The cloth’s are not yet placed where they belong, where should we put them? a) linen closet b) kitchen c) on the space heater d) fabric store e) bedroom Explanation: linen closets are typically used to store items such as cloths. Overall: 4.8 Well-written: Yes
Supports: a) Related: Yes
Factual: Yes
New info: 1.4
Unn. info: No
Contrastive: No
@@
Generated (worst) Sarah opened his chest and found a second heart. Her patient might not be what? a) istanbul b) animal c) antique shop d) human being e) bedroom Explanation: a human being is the only creature known to have two hearts. Overall: 1.2 Well-written: Yes
Supports: d) Related: Yes
Factual: No
New info: 0.8
Unn. info: No
Contrastive: No
@@
ECQA (best) Where would one find a captive monkey? a) rain forest b) tree tops c) toy store d) zoological gardens e) great outdoors Explanation: A monkey can be kept captive at a zoological garden as many animals are kept there for people to see. All the other options are incorrect as they do not have captive monkeys. Overall: 4.6 Well-written: Yes
Supports: c) Related: Yes
Factual: Yes
New info: 1.8
Unn. info: No
Contrastive: Yes
@@
ECQA (worst) Where are there lots of seats placed in rows surrounding a court? a) show b) auditorium c) movies d) soccer stadium e) hockey game Explanation: In auditorium, there are lots of seats placed in rows. Auditorium seats are surrounding a court. In show, there are no surrounding court placement of seats. Movies is not a place. In soccer stadium, seats are not placed in rows. In hockey game, there are no proper seats. Overall: 1.6 Well-written: No
Supports: b) Related: Yes
Factual: No
New info: 1.0
Unn. info: No
Contrastive: Yes
Table 4: Top-rated and bottom-rated examples of explanations for CommonsenseQA in our test set. The ratings on the right are aggregated from five crowdworkers’ annotations. Well-written ranges from $1.0$ to $5.0$, and New info. from $0.0$ to $3.0$.
COPA-SSE (best) The vehicles travelled slowly. What was the cause of this? a) It was foggy.
b) It was cloudy.
@@
Explanation: Fog is a thick mist which prevents you from driving quickly as you cannot see far in front of your vehicle. Driving slow through is motivated by the desire not to crash.
Overall: 4.8 Well-written: Yes
Supports: a) Related: Yes
Factual: Yes
New info: 2.0
Unn. info: No
Contrastive: No
@@
COPA-SSE (worst) The boy found the television show interesting. What happened as a result? a) He shut it off.
b) He recorded it.
Explanation: The result causes the record shut off.
Overall: 1.0 Well-written: No
Supports: None Related: Yes
Factual: N/A
New info: 0.4
Unn. info: Yes
Contrastive: No
COPA-SSE + Fluency fixed (best) The woman tolerated her friend’s difficult behavior. What was the cause of this? a) The woman knew her friend was going through a hard time.
b) The woman felt that her friend took advantage of her kindness.
@@
Explanation: Going through a hard time can cause people to act in a difficult way. A friend can be motivated to support their friend when they’re having a tough time.
Overall: 4.8 Well-written: Yes
Supports: a) Related: Yes
Factual: Yes
New info: 2.2
Unn. info: No
Contrastive: No
@@
COPA-SSE + Fluency fixed (worst) The shirt did not have a hole anymore. What was the cause of this? a) I patched the shirt.
b) I bleached the shirt.
Explanation: I desire to look good.
Overall: 1.0 Well-written: No
Supports: None Related: No
Factual: N/A
New info: 0.2
Unn. info: Yes
Contrastive: No
Generated (best) The boy was struggling to fall asleep. What was the cause of this? a) He turned on his nightlight.
b) He watched a horror movie.
@@
Explanation: Watching a horror movie can create fear and anxiety that can make it difficult for someone to fall asleep. The nightlight is not likely to be the cause since it is meant to help someone sleep.
Overall: 5.0 Well-written: Yes
Supports: b) Related: Yes
Factual: Yes
New info: 2.4
Unn. info: No
Contrastive: Yes
Generated (worst) The woman awakened. What was the cause of this? a) She took sleeping pills.
b) Her husband was snoring.

### v2

@@ Results.
Figure 5 shows Spearman’s rank correlation with the original gold ratings obtained by aggregating the labels of all five raters. Humans only (*H) denotes the correlation between human majority-voted labels only, and others when including the respective LLM as an additional rater (*H+LLM).
Figure 5: A comparison of Spearman’s rank correlation with the original gold labels when using fewer raters (*H) and when an LLM is added as an additional rater (*H+LLM). From left to right, the number of human raters decreases from four to one (randomly selected). All values are multiplied by 100 for legibility.
Each column cluster represents a different number of human raters from four in the leftmost to one in the rightmost. With four humans, the correlation between their majority-voted labels and the original gold labels was 0.91. Adding a model as a fifth vote raised this by only 0.004 points. With three humans, we observed a slight decrease in correlation when adding a model. With two or one human rater, the correlation increased by 0.019 and 0.016 points, respectively.
Overall, the results suggest that LLMs can be useful as additional raters when the number of human raters is less than three. However, even in the best case, the voted labels with an added LLM rater still scored lower than with an equivalent number of human raters (0.83 with 2H+LLM vs. 0.89 with 3H). When there is a high number of human raters, in this case three or more, the model’s inclusion as an additional rater does not improve the majority-voted labels’ alignment with the original gold labels, and may even harm it.
@@ 4 Discussion
We analyzed the alignment between human raters and LLMs in the context of evaluating the quality of explanations. The task is highly subjective, and some degree of variance is expected even between humans.
The best-performing models seemed to output labels that are within this variance, suggesting that they behave similarly to individual human raters. However, comparing its outputs to majority-voted labels revealed that their labels are not always in line with human majorities, indicating that they are not suitable as a complete replacement for human raters. In turn, when there were fewer human annotators, LLMs helped bring the majority-voted labels closer to the original gold labels. However, this improvement was not seen when there were already three or more human annotators (compared to the original five), suggesting that LLMs are only useful in extremely limited scenarios.
From this, we concluded that LLMs are not a reliable replacement for human raters, unless the use case does not require complete alignment. However, in this work, we operated under the assumption that majority-voted labels are the ground truth. In practice, this may not always be the case, and the reliability of LLMs may vary depending on what information is desired in a given scenario. This opens up a new avenue for future work, where LLMs can be applied in a more nuanced manner, considering the context of the task and the intended use of the labels.
Overall, as a balanced approach, we recommend using LLMs in configurations that end with targeted human involvement, such as extensively using LLMs in development or filtering stages and experts for final testing or evaluation. In future work, we plan to explore the behavior of LLMs more closely, particularly in identifying potential patterns in where the model diverged from humans, such as explanations with particular characteristics or in specific contexts. These insights could help us better understand the limitations of LLMs and how they can be used effectively in real-world applications or further improved.
@@ 5 Related Work
LLMs as data labelers, and more broadly humans-and-LLMs-in-the-loop settings, are an emerging direction in data collection and labeling. E.g., Wiegreffe et al. (2022) developed a predict-and-explain pipeline that combines GPT-3 with a supervised filter trained on binary acceptability judgments from humans. More recently, Chiang & Lee (2023) proposed using LLMs to evaluate text, closely in line with our work. They, however, found the models successful in their setting. In contrast, we focus on explanation rating with more complex fine-grained criteria, and closely scrutinize potential weaknesses; Previous works brought into question the reliability of LLMs’ predictions, especially in prompting setups. E.g., Webson & Pavlick (2022) found that instruction-tuned models often produced good predictions with irrelevant or misleading prompts, bringing into question their real “understanding” of the task. Others reported unreliability of LLMs as labelers in various settings (Albrecht et al., 2022; Shen et al., 2023; Hada et al., 2024, i.a.), a line of work which we join with findings in the context of explanation evaluation.
@@
For explanations in the form of textual justifications, previous works often defined their own criteria for evaluation. Automatic evaluation borrowed from machine learning and measures overlap with “gold standard” text using (a) word-overlap metrics, e.g., BLEU, METEOR and ROUGE; and (b) embedding-based metrics, e.g., BERTScore and BLEURT (Clinciu et al., 2021). In contrast, human-tagged measures are more diverse and explanation-specific. For example, Clinciu et al. (2021) measured Informativeness and Clarity; Wiegreffe et al. (2022), inspired by social sciences, measured Acceptability, Generality, Factuality, Grammar, New Info, Supports Label, and Amount of Information; while Aggarwal et al. (2021) defined the criteria of Refutation, Complete, Comprehensive, Minimal, and Coherent. A recent study instead focused on the utility of explanations, i.e., their helpfulness in answering a question from a human point of view (Joshi et al., 2023). In our work, we largely followed these existing criteria, with a modified version of the “supports” criterion to capture the consistency of a model’s predicted label with its justification.
@@
Explanations are often generated in predict-and-explain settings, where models provide justifications for their answers in QA-based benchmarks (e.g., Clinciu et al., 2021), or an elaborate-then-predict setting, where models output a prediction guided by its intermediate outputs such as knowledge statements or reasoning chains (e.g., Marasović et al., 2022; Wang et al., 2023). In this work, we evaluate the capabilities of LLMs to evaluate explanations in an explain-and-predict setting, where models provide justifications for their answers in commonsense reasoning benchmarks. The following sections provide background on each aspect.
@@ 6 Conclusions
Using a newly-built dataset of free-text explanations and crowdsourced aspect-wise quality ratings, we analyzed the viability of LLMs as explanation evaluators. Stronger LLMs increased or maintained inter-annotator agreement when replacing human ratings, and their ratings were moderately to highly correlated with human ratings, depending on the quality aspect. In a scenario where LLMs were used as additional raters instead of a complete replacement, they improved the outcome when there were only two human raters, but were neutral to detrimental when there are three or more human raters. We conclude that while LLMs can provide ratings moderately consistent with an average human rater, they are not yet reliable enough for complete replacement.
@@
Some of the commercial LLMs we used have built-in filters for potentially harmful content, which were triggered during several of our experiments. This data should not be used in any downstream applications without first controlling for potentially harmful samples.
@@ Appendix A Dataset Statistics
Criterion -1 0 1 2 3 4 5
Supports 11% 32% 31% 7% 9% 8%
Overall 16% 13% 28% 28% 12%
Well-written 27% 72%
Related 6% 93%
Factual 10% 4% 84%
New Info 30% 31% 35% 2%
Unnecessary Info 82% 17%
Contrastive 58% 41%
Table 3: Label distributions per criterion of majority-voted human ratings. -1 denotes "none" for supports and "N/A" for factual.
Table 2 shows a breakdown of the samples in the test set per source dataset as described in Section 2.2. Table 3 shows the label distributions of each criterion in our test set. Tables 5 and 6 show examples of best-rated and worst-rated explanations from each source dataset.
A.1 Dataset-Wise Average Ratings
Table 4 shows the mean ratings and standard deviations of majority-voted ratings per data subset, excluding the categorical aspect supports. Data labels are described in Section 2.1. All ratings are the higher the better, except for unnecessary information which is the lower the better.
Ovr. Well-wr. Rel. Fact. New i. Unn. i. Cntr.
Data source 1-5 0, 1 0, 1 -1, 0, 1 0-3 1, 0* 0, 1
CoS-E 1.89 ${}_{\text{0.95}}$ 0.34 ${}_{\text{0.47}}$ 0.77 ${}_{\text{0.42}}$ 0.23 ${}_{\text{0.95}}$ 0.32 ${}_{\text{0.51}}$ 0.41 ${}_{\text{0.49}}$ 0.01 ${}_{\text{0.09}}$
CoS-E + fl. fix 2.06 ${}_{\text{1.01}}$ 0.72 ${}_{\text{0.45}}$ 0.72 ${}_{\text{0.45}}$ 0.39 ${}_{\text{0.90}}$ 0.42 ${}_{\text{0.67}}$ 0.44 ${}_{\text{0.50}}$ 0.01 ${}_{\text{0.09}}$
CSQA generated 3.20 ${}_{\text{0.83}}$ 0.97 ${}_{\text{0.16}}$ 1.00 ${}_{\text{0.04}}$ 0.96 ${}_{\text{0.26}}$ 1.00 ${}_{\text{0.74}}$ 0.04 ${}_{\text{0.19}}$ 0.01 ${}_{\text{0.09}}$
ECQA 3.05 ${}_{\text{0.89}}$ 0.57 ${}_{\text{0.49}}$ 1.00 ${}_{\text{0.00}}$ 0.90 ${}_{\text{0.31}}$ 1.29 ${}_{\text{0.71}}$ 0.11 ${}_{\text{0.32}}$ 0.96 ${}_{\text{0.19}}$
COPA-SSE 2.38 ${}_{\text{1.13}}$ 0.45 ${}_{\text{0.50}}$ 0.91 ${}_{\text{0.28}}$ 0.58 ${}_{\text{0.74}}$ 0.84 ${}_{\text{0.80}}$ 0.33 ${}_{\text{0.47}}$ 0.02 ${}_{\text{0.13}}$
COPA-SSE + fl. fix 2.81 ${}_{\text{1.02}}$ 0.82 ${}_{\text{0.39}}$ 0.97 ${}_{\text{0.18}}$ 0.78 ${}_{\text{0.60}}$ 0.85 ${}_{\text{0.78}}$ 0.18 ${}_{\text{0.38}}$ 0.00 ${}_{\text{0.00}}$
BCOPA generated 4.21 ${}_{\text{0.79}}$ 1.00 ${}_{\text{0.04}}$ 1.00 ${}_{\text{0.00}}$ 0.97 ${}_{\text{0.20}}$ 1.75 ${}_{\text{0.63}}$ 0.00 ${}_{\text{0.06}}$ 0.95 ${}_{\text{0.21}}$
BCOPA crowdsourced 4.32 ${}_{\text{0.76}}$ 0.97 ${}_{\text{0.17}}$ 1.00 ${}_{\text{0.00}}$ 1.00 ${}_{\text{0.00}}$ 1.94 ${}_{\text{0.55}}$ 0.01 ${}_{\text{0.12}}$ 0.95 ${}_{\text{0.21}}$
All 3.07 ${}_{\text{1.27}}$ 0.72 ${}_{\text{0.45}}$ 0.93 ${}_{\text{0.25}}$ 0.75 ${}_{\text{0.63}}$ 1.11 ${}_{\text{0.87}}$ 0.17 ${}_{\text{0.38}}$ 0.42 ${}_{\text{0.49}}$
Table 4: Mean ratings and standard ${}_{\text{standard deviations}}$ per data subset. Higher is better for all criteria except for unnecessary information, marked with an asterisk (*), where lower is better.
The human raters seemed to find generated explanations to be most well-written on average, however, the higher quality human-written explanations (ECQA, BCOPA crowdsourced) had a higher amount of new information. The generated explanations, in turn, had the least amount of unnecessary information. Interestingly, even though they were not contrastive, the generated explanations for CommonsenseQA (CSQA generated) had a higher average overall rating than ECQA explanations which are explicitly contrastive.
A Random Forest Regressor, achieving a mean squared error of 0.37, deemed the most important predictive feature to be new information (58%), followed by factual (20%), unnecessary information (9%), well-written (7%), contrastive (4%), supports (2%), and related (0%). Note that for supports, we defined a binary feature of whether the label matches the answer label.
@@ Appendix C Preliminary Experiment: Prompting Strategies
We compared single and compound calling, where the former prompts the model for a single criterion at a time, and the latter prompts the model for all criteria at once. We also compared default, averaged, and verbatim prompt formats, corresponding to a simple prompt with the explanation and the rating criteria, a voting mechanism over several prompt variants, and an input identical to the human annotation guidelines, respectively. For single verbatim calls, only the relevant sections (guidelines and examples) for the target criterion were included. Default and averaged prompts were further compared in zero-shot and three-shot settings, where the latter contained the same examples as shown in the human guidelines. All models were most successful with verbatim compound prompting.
Appendix D Crowdsourced Explanations for BCOPA
Figure 7: Explanation writing AMT form.
@@ Appendix D Crowdsourced Explanations for BCOPA
The workers were explicitly instructed to write contrastive explanations which “focus on what the key difference between the options is, and how it leads to it being the correct choice in one case and not in the other.” This instruction was inspired by insights from Miller (2019). Figure 7 shows the crowdsourcing form. We collected two explanations per BCOPA question, totaling 3,000 explanations. Out of these, a random sample of 500 explanations was rated and included in our test set. Example explanations can be seen in Table 6.
@@
CoS-E (best) Which effect of stress could cause death if not treated immediately? a) age grieving person b) heart attacks c) depression d) hair loss e) headache Explanation: heart attacks can stop your heart from properly pumping blood, thus leading to death. Over.: 4 Well-wr.: yes
Supp.: b) Rel.: yes
Fact.: yes
New i.: ++
Unn. i.: no
Contr.: no
@@
CoS-E (worst) The college kids weren’t known for their apartment decor choices, they used a folding chair inside their what? a) bathroom b) closet c) school d) room e) church Explanation: after years of being held captive in an isolated Over.: 1 Well-wr.: no
Supp.: no Rel.: no
Fact.: n/a
New i.: -
Unn. i.: yes
Contr.: no
@@
CoS-E + fluency fixed (best) What kind of driving leads to accidents? a) stressful b) dangerous c) fun d) illegal e) deadly Explanation: Driving dangerously creates accidents. Over.: 4 Well-wr.: yes
Supp.: b) Rel.: yes
Fact.: yes
New i.: +
Unn. i.: no
Contr.: no
@@
CoS-E + fluency fixed (worst) What is a grumpy person likely to have? a) headache b) understand themselves c) compliments d) own house e) happiness Explanation: A peculiar person. Over.: 1 Well-wr.: no
Supp.: none Rel.: no
Fact.: n/a
New i.: -
Unn. i.: yes
Contr.: no
@@
Generated (best) The cloth’s are not yet placed where they belong, where should we put them? a) linen closet b) kitchen c) on the space heater d) fabric store e) bedroom Explanation: linen closets are typically used to store items such as cloths. Over.: 5 Well-wr.: yes
Supp.: a) Rel.: yes
Fact.: yes
New i.: ++
Unn. i.: no
Contr.: no
@@
Generated (worst) Sarah opened his chest and found a second heart. Her patient might not be what? a) istanbul b) animal c) antique shop d) human being e) bedroom Explanation: a human being is the only creature known to have two hearts. Over.: 1 Well-wr.: yes
Supp.: d) Rel.: yes
Fact.: no
New i.: -
Unn. i.: no
Contr.: no
@@
ECQA (best) Where would one find a captive monkey? a) rain forest b) tree tops c) toy store d) zoological gardens e) great outdoors Explanation: A monkey can be kept captive at a zoological garden as many animals are kept there for people to see. All the other options are incorrect as they do not have captive monkeys. Over.: 5 Well-wr.: yes
Supp.: d) Rel.: yes
Fact.: yes
New i.: ++
Unn. i.: no
Contr.: yes
@@
ECQA (worst) Where are there lots of seats placed in rows surrounding a court? a) show b) auditorium c) movies d) soccer stadium e) hockey game Explanation: In auditorium, there are lots of seats placed in rows. Auditorium seats are surrounding a court. In show, there are no surrounding court placement of seats. Movies is not a place. In soccer stadium, seats are not placed in rows. In hockey game, there are no proper seats. Over.: 2 Well-wr.: no
Supp.: b) Rel.: yes
Fact.: no
New i.: -
Unn. i.: no
Contr.: yes
Table 5: Top- and bottom-rated example explanations for CommonsenseQA in [Uncaptioned image]Acorn. New info. labels are denoted as follows: $+$ (ample), ++ (sufficient), + (some), - (none).
@@
COPA-SSE (best) The vehicles travelled slowly. Cause? a) It was foggy.     b) It was cloudy. Explanation: Fog is a thick mist which prevents you from driving quickly as you cannot see far in front of your vehicle. Driving slow through is motivated by the desire not to crash. Over.: 5 Well-wr.: yes
Supp.: a) Rel.: yes
Fact.: yes
New i.: ++
Unn. i.: no
Contr.: no
@@
COPA-SSE (worst) The boy found the television show interesting. Result? a) He shut it off.     b) He recorded it. Explanation: The result causes the record shut off. Over.: 1 Well-wr.: no
Supp.: none Rel.: yes
Fact.: n/a
New i.: -
Unn. i.: yes
Contr.: no
@@
COPA-SSE + Fluency fixed (best) The woman tolerated her friend’s difficult behavior. Cause? a) The woman knew her friend was going through a hard time.     b) The woman felt that her friend took advantage of her kindness. Explanation: Going through a hard time can cause people to act in a difficult way. A friend can be motivated to support their friend when they’re having a tough time. Over.: 5 Well-wr.: yes
Supp.: a) Rel.: yes
Fact.: yes
New i.: ++
Unn. i.: no
Contr.: no
@@
COPA-SSE + Fluency fixed (worst) The shirt did not have a hole anymore. Cause? a) I patched the shirt.     b) I bleached the shirt. Explanation: I desire to look good. Over.: 1 Well-wr.: no
Supp.: none Rel.: no
Fact.: n/a
New i.: -
Unn. i.: yes
Contr.: no
@@
Generated (best) The boy was struggling to fall asleep. Cause? a) He turned on his nightlight.     b) He watched a horror movie. Explanation: Watching a horror movie can create fear and anxiety that can make it difficult for someone to fall asleep. The nightlight is not likely to be the cause since it is meant to help someone sleep. Over.: 5 Well-wr.: yes
Supp.: b) Rel.: yes
Fact.: yes
New i.: ++
Unn. i.: no
Contr.: yes